Redes de difracción

Campo de interferencia producido por N rendijas de ancho a y espaciamento b:

I = I_{∘} [\frac{sin(\frac{π}{λ}a sin(θ))}{\frac{π}{λ}a sin(θ)} \frac{sin(\frac{π}{λ}Nb sin(θ))}{sin(\frac{π}{λ}b sin(θ))}]^{2}  (1)

Ejemplos para λ = 5000Å, a = 1mm, b = 5mm:

N = 2 N = 3 N = 100